# Monte Carlo Simulation Development and Implementation of the GiBUU Model for Neutrino Experiments

### v1

@@ I Introduction
In particular, accelerator-based neutrino experiments have significantly advanced our understanding of neutrinos and their properties. These experiments employ powerful particle accelerators to produce intense and well-characterized neutrino beams, allowing for detailed investigations of neutrino interactions, oscillations, and their potential implications for other physics beyond the Standard Model. Understanding neutrino oscillations is paramount for advancing our knowledge of the lepton sector and exploring new physics; for example, deviations from the observed oscillation patterns could indicate the presence of new neutrino flavors or interactions that go beyond the established framework of the Standard Model. Moreover, these experiments provide a unique avenue for probing the Universe’s asymmetry between matter and antimatter.
In this context, Monte Carlo simulations are essential for designing, analyzing, and interpreting experimental data. These simulations provide a means to model and predict neutrino interactions with target materials, detector responses, and backgrounds, thus enabling precise estimates of the expected event rates to extract the desired physics parameters, including corrections due to limited efficiencies and backgrounds. The importance of accurate models in Monte Carlo simulations stems from their significant impact on result extraction and the subsequent interpretation of physics outcomes.
This paper presents a Monte Carlo event generator based on the GiBUU (Giessen Boltzmann-Uehling-Uhlenbeck) model [3, 4] and specifically tailored for neutrino experiments. Within this work, the GiBUU model, renowned for its success in describing hadron/electron/photon/neutrino-nucleus interactions in various target materials, is converted into an event-by-event predictor. By incorporating the GiBUU model into a Monte Carlo simulation framework, we aim to provide a comprehensive tool for simulating and analyzing neutrino data events.
@@
The simulation leverages the GENIE [9, 10] geometry and flux drivers and produces the output event formats compatible with those from GENIE. By adopting this approach, we can run experiment-specific software to reconstruct the generated events in the same manner as we would with GENIE-generated events, eliminating the burden of developing additional software at each experiment to read GiBUU Monte Carlo outputs. The relevance of our Monte Carlo generator extends beyond the accurate modeling of neutrino-nucleus cross sections; it enables precise predictions of neutrino event rates, energy spectra, and particle kinematic distributions. These predictions are crucial for the interpretation of oscillation measurements and the search for new physics phenomena. The model accounts for various aspects, including final-state interactions and nuclear effects, and the generated events can quickly propagate to simulate the detector response using experiment-specific software, reproducing observed experimental signatures accurately using the same software setup as neutrinos generated by GENIE Monte Carlo.
Section II offers a brief description of the GiBUU model that is relevant for neutrino-nucleus scattering in the context of neutrino oscillations. Section III provides an overview and discussion of how the Monte Carlo event generator is produced. A brief discussion of the following steps towards calculating and implementing the systematic uncertainties associated with the GiBUU model is presented in Section IV, and the results of the work and final remarks are presented in Section V.
@@ II The GiBUU model and its relevance to neutrino interactions
The GiBUU (Giessen Boltzmann-Uehling-Uhlenbeck) [3, 4] model serves as a comprehensive theoretical framework designed to study interactions of hadrons, electrons, and neutrinos within various nuclear systems. Grounded in kinetic theory and effective quantum field theory principles, GiBUU incorporates nuclear transport theory and mean-field dynamics concepts to describe the evolution of hadronic systems by considering their interactions through scattering, decay, and production processes. The GiBUU model has found extensive use in a range of research areas, including nuclear reactions, heavy-ion collisions, and neutrino-nucleus interactions, the latter being the focus of this work. Its versatility and accuracy have made it valuable in interpreting experimental data and understanding the underlying dynamics of nuclear systems.
@@
A essential feature of the GiBUU model is its remarkable accuracy in replicating various experimental datasets without the need for tuning to neutrino scattering data, in which several assumptions of the initial neutrino would have been to be made. To develop some of the model parametrizations, the GiBUU model contains tunings to electron and hadron experimental scattering data, where precise knowledge of the incoming particle energy is available. GiBUU has been extensively tested and validated against various measurements, demonstrating its stability and robustness in describing neutrino-nucleus interactions, including argon targets [11]. This unique feature of the GiBUU model highlights its inherent ability to capture the intricacies of neutrino interactions without relying on specific adjustments or tuning procedures to match neutrino experimental results. Therefore, using the GiBUU model for calculating the efficiency, the background description, and other model-dependent measurement corrections is expected to reduce associated systematic uncertainties. Furthermore, the advantage of not relying on neutrino data tuning allows for more robust handling of the related uncertainties, as comparisons of the model with neutrino scattering data can help estimate some of the model’s uncertainties. By leveraging the inherent accuracy and stability of the GiBUU model, the systematic uncertainties associated with neutrino experiments can be effectively addressed and minimized, enhancing the reliability and precision of the results obtained.
GiBUU treats the neutrino-nucleus interaction using weak charged current (CC) and neutral current (NC) interaction processes. It considers the quasi-elastic scattering (QE), resonant and non-resonant production mechanisms, deep inelastic scattering, and meson-exchange current processes that play significant roles in neutrino interactions.
@@
The GiBUU model intricately incorporates nuclear effects via a comprehensive treatment of Final State Interactions (FSI) with nucleon-level interactions. FSI accounts for the intricate interplay between the produced hadrons and the encompassing nuclear medium, including rescattering absorption and re-emission processes. To meticulously capture these interactions with nuclei, the model integrates nuclear structure information, maintains a coherent nuclear potential description, and employs the transport model (the BUU equation). This approach ensures a precise treatment of nuclear effects and elucidates the nuanced influence of a nucleus on these interactions. In GiBUU, this facet proves indispensable for the comprehension of hadronic observables arising from the presence of nuclear matter. This aspect is crucial for neutrino experiments analyzing exclusive neutrino scattering topologies, in which the description of each of the particles produced leaving the nucleus plays an essential role in neutrino energy reconstruction.
@@ III Monte Carlo Simulation Framework for Neutrino Experiments
One of the most common neutrino Monte Carlo event generators, GENIE [9, 10], has emerged as an essential tool for simulating neutrino interactions in various experiments, and in particular for liquid argon experiments at Fermilab, such as MicroBooNE [14], ICARUS [7], SBND [6] and DUNE [15] on which coordinated efforts for development, tuning, and uncertainty estimation of the GENIE generator are already formed. By generating simulated events that mimic real neutrino interactions, GENIE facilitates interpreting data obtained from neutrino experiments and aids in developing and optimizing experimental designs. The widespread utilization of the GENIE event generator in numerous neutrino experiments establishes its critical role. Consequently, it is unsurprising that any subsequent software infrastructure development encompassing detector simulation and event reconstruction is inherently compatible with the output format of GENIE for many experiments and, in particular, for USA-based experiments. Hence, within this work, the simulation of neutrino events using the GIBUU model, considering a neutrino beam and detector geometry (LArBox), utilizes the GENIE flux and geometry drivers. Furthermore, the resulting generated GiBUU events are formatted in accordance with GENIE’s format to ensure seamless integration and compatibility.
Our simulation consists of two distinct steps. Firstly, a significant number of GiBUU events are simulated utilizing the original GiBUU model framework that the Giessen group developed. We produced $1.3\times 10^{10}$ events aiming to generate extensive libraries of neutrino events. This approach ensures event uniqueness and minimizes repetitions for each neutrino flavor and interaction type based on a given neutrino flux. The original GiBUU simulation provides cross-sections, including interaction amplitudes, some carrying negative weights, signifying their contribution as destructive interference terms in the overall cross-section calculation. These negative weights arise from non-resonant pion background processes, particularly relevant at higher energies such as those expected at the DUNE experiment. Each entry of the libraries stores the kinematic information of the incoming neutrino and produced particle kinematics, its GiBUU weight, and the event generator reaction code, which denotes the initial neutrino-nucleus interaction mode. Two libraries are created; one contains events with positive weights and the other with negative weights. In the latter case, the negative weights are converted to positive by taking their absolute values.
@@
Secondly, two independent productions are made, one per library set, using an acceptance-rejection method, event by event, to select the input from the library. The weight distribution from narrow neutrino energy in the library is used as probability distribution function input to the method, like the sample shown in Fig. 1. This visualization also provides insights into the scale of the generated libraries and the corresponding weight yields within a narrow energy range. Our strategy is to run any high-level analysis with the two productions and to calculate the desired final kinematic distribution by subtracting them. Furthermore, in addition to building a comprehensive statistical library of GiBUU events, we apply interaction vertex rotation techniques to avoid event duplication. Figure 2 shows the neutrino energy distribution for the two productions in the LArBox geometry. The ”negative weights” curve is significantly lower than the ”positive weights” at these energies.
@@
Figure 1: Distribution of weights in the GiBUU library generated for this study for muon neutrino charged-current interaction for neutrino energy in $0.600-0.601$ GeV.
@@
Figure 2: Neutrino energy distribution in the LArBox geometry for the ”negative” and ”positive” weights.
@@
Our code is a customized version of the GENIE evtLib [10], which diverges from the original in several aspects. Notably, our customized evtLib can handle weighted events and crucially preserves the event identity, retaining the information related to the event generator reaction code. Subsequently, these events are incorporated into the LArSoft framework as an external library. The flux and detector geometry drivers developed within the GENIE framework are employed during this stage. The same methodology could be used for other cross-section models not yet available as event generators, such as some Beyond the Standard Model theory simulations. Our simulation study employed the described technique to simulate GiBUU events within the LArBox geometry, utilizing the BNB (Booster Neutrino Beam) neutrino beam [8] prediction.
@@
Following this, we conducted several validation procedures to check the authenticity of the produced events by comparing them to the original GiBUU cross-section library, ensuring their conformity to the original GiBUU model. Figure 3 shows two of these studies looking at $\nu_{\mu}$ charged-current ($\nu_{\mu}$ CC) events with respect to the neutrino energy: the average hadron multiplicity and the ratio between events with a MEC-enhanced selection (one muon, two protons, and zero pions in the final state) and a QE-enhanced selection (one muon, one proton, and zero pions in the final state). Besides statistical fluctuations, especially at larger energies (where the BNB flux at LArBox is negligible) and minor differences due to the geometrical effects, the validation was successfully completed. The validation tests confirmed the GiBUU model’s transformation into a Monte Carlo event generator incorporating the appropriate flux and geometry settings.
@@
Figure 3: Validation plots comparing the GiBUU cross-section library vs. GiBUU events generated in the LArBox geometry. The plot on the top shows the average hadron multiplicity per $\nu_{\mu}$ CC interactions with respect to the neutrino energy. The bottom plot calculates the ratio between a selection that enhances MEC (one muon, two protons, and zero pions) and a selection that enhances QE interactions (one muon, one proton, and zero pions).
IV Unaddressed Developments in Systematical Uncertainty Estimation for GiBUU

### v2

@@ I Introduction
In particular, accelerator-based neutrino experiments have significantly advanced our understanding of neutrinos and their properties. These experiments employ powerful particle accelerators to produce intense neutrino beams, allowing for detailed investigations of neutrino interactions, oscillations, and their potential implications for other physics beyond the Standard Model. Understanding neutrino oscillations is paramount for advancing our knowledge of the lepton sector and exploring new physics; for example, deviations from the observed oscillation patterns could indicate the presence of new neutrino flavors or interactions that go beyond the established framework of the Standard Model. Moreover, these experiments provide a unique avenue for probing the Universe’s asymmetry between matter and antimatter.
In this context, Monte Carlo simulations are essential for designing, analyzing, and interpreting experimental data. These simulations provide a means to model and predict neutrino interactions with target materials, detector responses, and backgrounds, thus enabling precise estimates of the expected event rates to extract the desired physics parameters, including corrections due to limited efficiencies and backgrounds. The importance of accurate models in Monte Carlo simulations stems from their significant impact on result extraction and the subsequent interpretation of physics outcomes. The relevance of our Monte Carlo generator extends beyond the accurate modeling of neutrino-nucleus cross sections; it enables precise predictions of neutrino event rates, energy spectra, and particle kinematic distributions. These predictions are crucial for the interpretation of oscillation measurements and the search for new physics phenomena.
This paper presents a Monte Carlo event generator based on the GiBUU (Giessen Boltzmann-Uehling-Uhlenbeck) model [3, 4] and specifically tailored for neutrino experiments. The output of GiBUU is a weighted event simulation, which adds difficulty to its implementation as an event simulation within a detector geometry. Within this work, the GiBUU model, renowned for its success in describing hadron/electron/photon/neutrino-nucleus interactions in various target materials, is converted into an event-by-event predictor, or in other words, an unweighted event generator. By incorporating the GiBUU model into a Monte Carlo simulation framework, we aim to provide a comprehensive tool for simulating and analyzing neutrino data events. An important aspect of neutrino oscillations is reconstructing the neutrino energy, mostly relying on the final state particles observed from the neutrino interaction. The model approach adopted by GiBUU with its nuclear transport model, different from GENIE and other neutrino event generators, can offer important insights to understand those final state particles and improve our accuracy to assess uncertainties in corrections due to inefficiencies and backgrounds. GiBUU cannot account for detector geometry, so it was unsuitable to be implemented as an event generator for experiments.
The simulation leverages the GENIE [5, 6] geometry and flux drivers and produces the output event formats compatible with those from GENIE. By adopting this approach, we can run experiment-specific software to reconstruct the generated events in the same manner as we would with GENIE-generated events, eliminating the burden of developing additional software at each experiment to read GiBUU Monte Carlo outputs. The generated events can be efficiently propagated to simulate detector responses using experiment-specific software. This process accurately reproduces observed experimental signatures with the same software setup used for neutrinos generated by the GENIE Monte Carlo. Notably, the developed method also applies to other event generators, which could be especially beneficial for those producing weighted events, a common feature in many Beyond the Standard Model simulations.
@@
Section II offers a brief description of the GiBUU model that is relevant for neutrino-nucleus scattering in the context of neutrino oscillations. Section III provides an overview and discussion of how the Monte Carlo event generator is produced. A brief discussion of the following steps towards calculating and implementing the systematic uncertainties associated with the GiBUU model is presented in Section IV, and the results of the work, including comparisons to GENIE, and final remarks are presented in Section V.
@@ II The GiBUU Model and Its Relevance to Neutrino Interactions
The GiBUU [3, 4] model serves as a comprehensive theoretical framework designed to study interactions of hadrons, electrons, and neutrinos within various nuclear systems. Grounded in kinetic theory and effective quantum field theory principles, GiBUU incorporates nuclear transport theory and mean-field dynamics concepts to describe the evolution of hadronic systems by considering their interactions through scattering, decay, and production processes. The GiBUU model has been extensively used in a range of research areas, including nuclear reactions, heavy-ion collisions, and neutrino-nucleus interactions, the latter being the focus of this work. Its versatility and accuracy have made it valuable in interpreting experimental data and understanding the underlying dynamics of nuclear systems.
@@
An essential feature of the GiBUU model is its remarkable accuracy in replicating various experimental datasets without the need for tuning to neutrino scattering data, in which several assumptions of the initial neutrino would have been to be made. To develop some of the model parametrizations, the GiBUU model is informed by comparisons to electron and hadron experimental scattering data, where precise knowledge of the incoming particle energy is available [11]. GiBUU has been extensively tested and validated against various measurements, demonstrating its stability and robustness in describing neutrino-nucleus interactions, including argon targets [12]. This unique feature of the GiBUU model highlights its inherent ability to capture the intricacies of neutrino interactions without relying on specific adjustments or tuning procedures to match neutrino experimental results. Therefore, using the GiBUU model to calculate efficiency, background description, and other model-dependent measurement corrections can provide important insights to understand uncertainties in the physics models and prevent shortcomings. Understanding how the uncertainties are associated with the model, without relying on neutrino data tuning, enables a more robust handling of related uncertainties. By leveraging the inherent accuracy and stability of the GiBUU model, the systematic uncertainties associated with neutrino experiments can be effectively addressed and can help us to advance the efforts to minimize those uncertainties, enhancing the reliability and precision of the results obtained.
GiBUU treats the neutrino-nucleus interaction using weak charged current (CC) and neutral current (NC) interaction processes. It considers the quasi-elastic scattering (QE), resonant and non-resonant production mechanisms, deep inelastic scattering (DIS), and meson-exchange current processes that play significant roles in neutrino interactions. The nuclear ground model uses a Woods-Saxon density distribution while ensuring a constant Fermi-energy. Some of the neutrino-nucleus interaction mechanisms, such as 1-pion background and 2-pion background, are modeled using MAID [13] predictions, while PYTHIA [14] predictions are applied for DIS. Quasi-elastic and resonant processes are treated simultaneously when calculating lepton-nucleon interaction cross-sections for a single-hadron final state as quasi-elastic scattering and resonance excitation. Other processes, such as lepton-nucleon interaction cross sections for a two-hadron final state, contribute to $N\pi$ backgrounds. Meson exchange current interactions are implemented to CC and NC events using the same model from Bosted and Christy [15]. For a more detailed description of the interaction models, see [16].
@@
The GiBUU model incorporates nuclear effects through a comprehensive treatment of Final State Interactions (FSI) with nucleon-level interactions. FSI accounts for the intricate interplay between the produced hadrons and the encompassing nuclear medium, including rescattering absorption and re-emission processes. To meticulously capture these interactions with nuclei, the model integrates nuclear structure information, maintains a coherent nuclear potential description, and employs the transport model (the BUU equation). This approach ensures a precise treatment of nuclear effects and elucidates the nuanced influence of a nucleus on these interactions. In GiBUU, this facet proves indispensable for the comprehension of hadronic observables arising from the presence of nuclear matter. The FSI model in GiBUU is more sophisticated than in other event generators, and systematic uncertainties due to FSI may differ from the approach followed by other event generators. The work to understand these uncertainties and how to implement them is ongoing. This aspect is crucial for neutrino experiments analyzing exclusive neutrino scattering topologies, in which the description of each produced particle leaving the nucleus plays an essential role in neutrino energy reconstruction.
@@ III Monte Carlo Simulation Framework for Neutrino Experiments
One of the most common neutrino Monte Carlo event generators, GENIE [5, 6], has emerged as an essential tool for simulating neutrino interactions in various experiments, and in particular for liquid argon experiments at Fermilab, such as MicroBooNE [19], ICARUS [9], SBND [8] and DUNE [20] in which coordinated efforts for development, tuning, and uncertainty estimation of the GENIE generator are already formed. By generating simulated events that mimic real neutrino interactions, GENIE facilitates interpreting data obtained from neutrino experiments and aids in developing and optimizing experimental designs. The widespread utilization of the GENIE event generator in numerous neutrino experiments establishes its critical role. Consequently, it is unsurprising that any subsequent software infrastructure development encompassing detector simulation and event reconstruction is inherently compatible with the output format of GENIE for many experiments and, in particular, for USA-based experiments. Hence, within this work, the simulation of neutrino events using the GIBUU model, considering a neutrino beam and detector geometry (LArBox), utilizes the GENIE flux and geometry drivers. Furthermore, the resulting generated GiBUU events are formatted following GENIE’s format to ensure seamless integration and compatibility.
Our simulation consists of two distinct steps.
Firstly, a significant number of GiBUU events are simulated utilizing the original GiBUU model framework that the Giessen group developed. We produced $1.3\times 10^{10}$ events aiming to generate extensive libraries of neutrino events. This approach ensures event uniqueness and minimizes repetitions for each neutrino flavor and interaction type based on a given neutrino flux.
The original GiBUU simulation provides cross-sections, including the squares of the interaction amplitudes, some carrying negative weights, signifying their contribution as destructive interference terms in the overall cross-section calculation. These negative weights arise from non-resonant pion background processes, particularly relevant at higher energies such as those expected at the DUNE experiment. For example, for pion production, the sum of the square of the background amplitude and the interference term of resonance and background amplitudes can be negative. This contribution is the event type that describes the $1\pi^{+}$ background plus interference.
Each entry of the libraries stores the kinematic information of the incoming neutrino and produced particle kinematics, its GiBUU weight, and the event generator reaction code, which denotes the initial neutrino-nucleus interaction mode. Two libraries are created; one contains events with positive weights and the other with negative weights. In the latter case, the negative weights are converted to positive by taking their absolute values.
@@
Secondly, two independent productions are made, one per library set, using an acceptance-rejection method, event by event, to select the input from the library. The weight distribution from narrow neutrino energy in the library is used as probability distribution function input to the method, like the sample shown in Fig. 1. This visualization also provides insights into the scale of the generated libraries and the corresponding weight yields within a narrow energy range. Our strategy is to run any high-level analysis with the two productions and to calculate the desired final kinematic distribution by subtracting them.
Furthermore, in addition to building a comprehensive statistical library of GiBUU events, we apply interaction vertex rotation techniques to avoid event duplication. Figure 2 shows the neutrino energy distributions for muon-neutrino charged-current ($\nu_{\mu}CC$) events for both productions in the LArBox geometry. The ”negative weights” curve is significantly lower than the ”positive weights” at these energies.
@@
Figure 1: Distribution of weights generated in the GiBUU library for this study, illustrating muon neutrino charged-current interactions at a neutrino energy range of $0.6-0.601$ GeV.
@@
Figure 2: Neutrino energy distribution in the LArBox geometry for muon neutrino charged-current interactions, showing both ”negative” and ”positive” weight contributions.
@@
Our code is a customized version of the GENIE evtLib [6], which diverges from the original in several aspects. Notably, our customized evtLib can handle weighted events and crucially preserves the event identity, retaining the information related to the event generator reaction code (QE, resonant, 2p2h,..), run, and event ID as a unique identifier per event, as well as neutrino energy, weight, and an array of final state particles with their kinematics. Subsequently, these events are incorporated into the LArSoft framework as an external library. During this stage, the flux and detector geometry drivers developed within the GENIE framework are employed to place the vertices for the pre-calculated cross sections from the GiBUU libraries. The same methodology could be used for other cross-section models not yet available as event generators, such as some Beyond the Standard Model theory simulations.
Our simulation study employed the described technique to simulate GiBUU events within the LArBox geometry, utilizing the BNB (Booster Neutrino Beam) neutrino beam [10] prediction. For the studies presented in this paper, we used the GHEP event format. However, it is important to note that other event formats available in GENIE can be used, such as GST, rootracker, and other future event generator formats, such as NuHEPMC, that may be included in GENIE. This facilitates our software implementation in any experiment already using GENIE.
Following this, we conducted several validation procedures on the implementation, including verifying the accuracy of the event rate calculation. Figure 3 shows the ratios of inclusive charged-current cross sections (GiBUU over GENIE) as a function of neutrino energy, compared with ratios of generated events with equivalent neutrino flux exposure, demonstrating that both ratios are identical.
Figure 3: Validation plot comparing the ratios of inclusive charged-current cross sections (GiBUU over GENIE, shown with solid lines) compared with ratios of generated events (shown with dotted lines) using equivalent neutrino flux exposure.
Additional validations were performed to verify the authenticity of the produced events by comparing them to the original GiBUU cross-section library, ensuring their conformity to the original GiBUU model. Figure 4 shows two of these studies looking at $\nu_{\mu}CC$ events with respect to the neutrino energy: the average hadron multiplicity and the ratio between events with a MEC-enhanced selection (one muon, two protons, and zero pions in the final state) and a QE-enhanced selection (one muon, one proton, and zero pions in the final state). Within the different tests, we studied the topological description of the simulated neutrino events and the contributing neutrino-nucleus interaction types. The interaction types, such as QE, 2p2h, resonant, and DIS, provide essential information to understand the details of the modeling and comparisons to other generators.
@@
Besides statistical fluctuations, especially at larger energies (where the BNB flux at LArBox is negligible) and minor differences due to the geometrical effects, the validation was successfully completed. The validation tests confirmed the GiBUU model’s transformation into a Monte Carlo event generator incorporating the appropriate flux and geometry settings.
@@
Figure 4: Validation plots comparing the GiBUU cross-section library with GiBUU-generated events in the LArBox geometry as a function of neutrino energy. Top plot: Average hadron multiplicity per $\nu_{\mu}$ CC interaction. Bottom plot: Ratio of a selection enhancing MEC (one muon, two protons, zero pions) to a selection enhancing QE interactions (one muon, one proton, zero pions).
However, a topological event description, such as classifying events according to their final state particles, is also required to better understand contributions from final-state interactions (FSI). Comparing the final-state particle contribution per event to measured data is the most effective approach, as only these final-state particles can be detected in experiments. Furthermore, different neutrino-nucleus interactions and FSI can lead to similar observed topologies. It is also important to note that in GiBUU, final state protons below 9.38 MeV in kinetic energy in argon interactions should not be considered final state particles as they may still be bound to the nucleus. In the studies presented in Section V, we incorporate a minimum kinetic energy threshold at each hadron to describe each topology to anticipate the detectability capabilities of current LArTPCs and avoid bound protons with low kinetic energy.
IV Ongoing Developments in Systematic Uncertainty Estimation for GiBUU